# Multi-scale analysis of flow over heterogeneous urban environments

### v1

@@ 2.3 Triple decomposition
We define the Reynolds average of $\varphi$ as an ensemble average:
$\overline{\varphi}(\boldsymbol{x},t)=\sum_{i=1}^{N}\varphi^{(i)}(\boldsymbol{x},t)\,.$ (10)
Deviations from the average value $\overline{\varphi}$ are denoted $\varphi^{\prime}(\boldsymbol{x},t)=\varphi(\boldsymbol{x},t)-\overline{\varphi}(\boldsymbol{x},t)$. For statistically steady problems, time-averaging can be used in lieu of an ensemble average. In the current paper, we will assume that Reynolds-averaging has been performed before spatial averaging, which avoids some of the complexities of simultaneous Reynolds and spatial averaging (Schmid et al. 2019). We thus decompose the average velocity into the local spatial average and its deviation from the average as $\overline{\varphi}=\left\langle{\smash{\overline{\varphi}}}\right\rangle{{}^{f}}+\overline{\varphi}^{\prime\prime}$. This decomposition is only defined in the fluid domain, and the intrinsic average is used since this represents the average value inside the fluid domain. Substitution into (10) results in the triple decomposition
@@
For products, e.g. of the vertical scalar flux $w\varphi$, this implies
$\begin{split}\left\langle{\smash{\overline{w\varphi}}}\right\rangle&=\left\langle{\smash{\overline{w}\,\overline{\varphi}}}\right\rangle+\left\langle{\smash{\overline{w^{\prime}\varphi^{\prime}}}}\right\rangle\\ &=\left\langle{\smash{\left\langle{\smash{\overline{w}}}\right\rangle{{}^{f}}\left\langle{\smash{\overline{\varphi}}}\right\rangle{{}^{f}}}}\right\rangle+\left\langle{\smash{\left\langle{\smash{\overline{w}}}\right\rangle{{}^{f}}\overline{\varphi}^{\prime\prime}}}\right\rangle+\left\langle{\smash{\overline{w}^{\prime\prime}\left\langle{\smash{\overline{\varphi}}}\right\rangle{{}^{f}}}}\right\rangle+\left\langle{\smash{\overline{w}^{\prime\prime}\overline{\varphi}^{\prime\prime}}}\right\rangle+\left\langle{\smash{\overline{w^{\prime}\varphi^{\prime}}}}\right\rangle\,.\\ \end{split}$ (12)
None of these terms can be expected to be zero a priori, but note that for periodic domains $\left\langle{\smash{\overline{w}}}\right\rangle{{}^{f}}=0$ and $\left\langle{\smash{\overline{w}^{\prime\prime}\left\langle{\smash{\overline{\varphi}}}\right\rangle{{}^{f}}}}\right\rangle=0$ for $L\gg\ell$ (the latter term being zero since $\left\langle{\smash{\overline{\varphi}}}\right\rangle{{}^{f}}$ is homogeneous in the horizontal directions in this limit).
@@ 2.4 Reynolds- and area-averaged streamwise momentum equations
where $\delta_{ij}$ is the Kronecker delta. Superficial averaging and application of the spatial averaging theorem (7) results in
@@
$\left\langle{\smash{f_{D}}}\right\rangle=-\oint_{\partial\Omega_{f}}\mathcal{A}\frac{F_{i}N_{i}}{|\boldsymbol{N}_{\perp}|}\textnormal{d}s=-\oint_{\partial\Omega_{f}}\mathcal{A}\,\left(\overline{p}\frac{N_{x}}{|\boldsymbol{N}_{\perp}|}-\nu\frac{\partial\overline{u}}{\partial x_{j}}\frac{N_{j}}{|\boldsymbol{N}_{\perp}|}\right)\textnormal{d}s\,,$ (16)
since mean and fluctuating components are zero on the solid boundaries.
Considering the fact that NWP models take into account the net effect of buildings without actually resolving them, the governing equations are posed in terms of superficial quantities. Recasting (15) in the standard form for the streamwise momentum equation, we thus obtain:
@@
$\left\langle{\smash{\tau_{i}}}\right\rangle=\left\langle{\smash{\overline{u}_{i}}}\right\rangle\left\langle{\smash{\overline{u}}}\right\rangle-\left\langle{\smash{\overline{u}_{i}\overline{u}}}\right\rangle-\left\langle{\smash{\overline{u^{\prime}_{i}u^{\prime}}}}\right\rangle$ (18)
is the unresolved kinematic stress term. Note that we have not decomposed the term $\left\langle{\smash{\overline{u}_{i}\overline{u}}}\right\rangle$ into its components and that the viscous flux is negligible in magnitude compared to the other fluxes. Equation (17) highlights the two terms that need to be closed in order to represent the effect of buildings in a non-building resolving NWP model: 1) the distributed drag contribution $f_{D}$; and 2) the unresolved kinematic stress $\tau_{i}$.
@@ 2.5 Statistically steady flow with periodic boundary conditions
The case considered in this paper is a turbulent urban boundary layer flow with periodic boundary conditions and a free-slip top boundary at $z=h$ in a statistically steady state. In this situation, (17) becomes
@@
which is not much of a simplification. No term except the time-derivative can be assumed zero for arbitrary averaging length $L$. However, if the averaging length scale is very large, the problem will become homogeneous in $x$ and $y$ and all the horizontal derivatives will vanish. With this in mind, it is useful to rewrite the equation above as
$\frac{\partial\left\langle{\smash{\overline{u}}}\right\rangle\left\langle{\smash{\overline{u}}}\right\rangle}{\partial x}+\frac{\partial\left\langle{\smash{\overline{v}}}\right\rangle\left\langle{\smash{\overline{u}}}\right\rangle}{\partial y}+\frac{\partial\left\langle{\smash{\overline{p}}}\right\rangle}{\partial x}-\frac{\partial\left\langle{\smash{\tau_{x}}}\right\rangle}{\partial x}-\frac{\partial\left\langle{\smash{\tau_{y}}}\right\rangle}{\partial y}=\left\langle{\smash{f}}\right\rangle-\left\langle{\smash{f_{D}}}\right\rangle+\frac{\partial\left\langle{\smash{\tau_{z}}}\right\rangle}{\partial z}.$ (20)
Here, the horizontal terms on the left-hand side of the will vanish if the problem is homogeneous, leaving the terms on the right-hand side of the equation to balance. The equation above indicates that an imbalance in the terms on the right-hand side implies inhomogeneity in $x$ and $y$ via horizontal terms on the left-hand side and vice versa.
@@
$\left\langle{\smash{\tau_{z}}}\right\rangle_{\infty}=-\left\langle{\smash{\overline{w}\,\overline{u}}}\right\rangle_{\infty}-\left\langle{\smash{\overline{w^{\prime}u^{\prime}}}}\right\rangle_{\infty}=-\left\langle{\smash{\overline{w}^{\prime\prime}\overline{u}^{\prime\prime}}}\right\rangle_{\infty}-\left\langle{\smash{\overline{w^{\prime}u^{\prime}}}}\right\rangle_{\infty}\,.$ (22)
The subscript on the superficial average indicates the averaging length $L$, in this case $L/\ell\gg 1$ which we denote as $L=\infty$. In the limit of $L\gg\ell$, the quantity $\left\langle{\smash{\varphi}}\right\rangle_{\infty}$ is identical to the conventional plane-average $\left\langle{\smash{\varphi}}\right\rangle_{xy}$, which is typically used in computational studies (Coceal et al. 2006; Xie et al. 2008; Giometto et al. 2016; Nazarian et al. 2020; Sützl et al. 2020). The superficial plane average of $\varphi$ is defined as
$\left\langle{\smash{\varphi}}\right\rangle_{xy}(z,t)=\frac{1}{A_{T}}\int_{\Omega_{f}(z)}\varphi(\boldsymbol{x},t)\,\textnormal{d}\boldsymbol{x}_{\perp}\,,$ (23)
@@
Integrating (21) from height $z$ to the domain top at height $h$ results in the following cumulative stress relation (Sützl et al. 2020):
@@
$\tau_{f;L}(\boldsymbol{x}_{\perp},z)$ $=\int_{z}^{h}\left\langle{\smash{f}}\right\rangle_{L}(\boldsymbol{x}_{\perp},z^{\prime})\textnormal{d}z^{\prime}\,,$
$\tau_{D;L}(\boldsymbol{x}_{\perp},z)$ $=\int_{z}^{h}\left\langle{\smash{f_{D}}}\right\rangle_{L}(\boldsymbol{x}_{\perp},z^{\prime})\textnormal{d}z^{\prime}\,,$
@@
In order to keep consistency in notation with earlier work, we will use that in the absence of a specific filter size in the subscript of a quantity, it is implied that $L=\infty$, e.g. $\tau_{f}\equiv\tau_{f;\infty}$; and we will use the conventional plane average notation $\left\langle{\smash{\varphi}}\right\rangle_{xy}$ instead of its equivalent $\left\langle{\smash{\varphi}}\right\rangle_{\infty}$. With this in mind, the cumulative stress relation (24) is simply
@@ 3.1 Case setup
In order to investigate multi-scale aspects of urban flows, we consider an idealised quasi-realistic urban morphology that has a staggered arrangement of cuboidal buildings of breadth and width $W=24\,\textnormal{m}$, whose height $H$ is random (Figure 2). This case is inspired by Xie et al. (2008), but with a much larger simulation domain of $L_{x}\times L_{y}\times L_{z}=1\,536\,\textnormal{m}\times 768\,\textnormal{m}\times 512\,\textnormal{m}$, and using a height distribution representative for London (Sützl et al. 2021) rather than a Gaussian height distribution.
@@
The mean building height is set to $h_{m}=30\,\textnormal{m}$ and the ratio is set to $r=2.5$, both of which are close to the average value of London reported by Sützl et al. (2021). Thus the maximum building height is $h_{\max}=75\,\textnormal{m}$ and $\alpha=2.61$. At $z=0$, the total building width $L_{b}(z;\alpha)=N_{b}W$, which implies that the frontal area can be determined from (26) as
@@
In order to assign heights to each cuboidal building, note that (26) divided by $W$ can be inverted to show the height $z$ as a function of the number of buildings $L_{b}(z)/W$, which represents the distribution of heights as a function of the number of buildings (figure 2(a), solid line). The building heights can be determined by uniformly sampling this function for each $i=1,2,\ldots,N_{b}$. The first X buildings, where the function is not defined, are set to $h_{\max}$, and the minimum building height is set to $10$ m. As a final step, the heights are randomly distributed using a uniform distribution to all $512$ buildings as shown in the morphology in figure 2(b). The total building width $L_{b}$ is shown in the figure 6(a) against the height, above $z=10$ m which is the minimum building height, the building width $L_{b}$, also indicating the distributive frontal area smoothly decrease to zero to the top of the canopy region. Due to the minimum building height setting, the actual frontal area index is $\lambda_{f}=0.29$, slightly greater than previously designed.
@@
The urban flows over this morphology are performed by the Large-eddy simulation using our open-source code uDALES (Owens et al. 2024), presenting the solid boundary using the immersed boundary method (IBM). The dynamics near the boundary are parametrised by the logarithmic wall functions (Uno et al. 1995; Suter et al. 2022). The turbulent eddy viscosity is calculated following the Vreman (2004) subgrid model. The code employs a second-order central difference scheme on a staggered Arakawa C-grid for spatial discretization and an explicit third-order Runge-Kutta scheme for time integration.
[…]
3.2 Planar-averaging implementation details
@@
Efficient computation of the convolution integral (2) is crucial to carry out a multi-scale analysis. In the spatial domain, the convolution integral needs to be evaluated $N_{x}\times N_{y}\times N_{z}$ times for a single averaging length $L$ which is computationally very expensive, particularly for high-resolution simulations. Therefore, we perform the convolution in the Fourier space by making use of the convolution theorem which states that a convolution becomes a multiplication in Fourier space (Bracewell 2000). We thus compute the horizontal 2-D FFT of $I_{f}\varphi$ and the filter $\mathcal{A}$, multiply them, and then transform the product back to physical space using an inverse FFT, obtaining the filtered results for all points in the plane simultaneously. This procedure is repeated $N_{z}$ times to obtain the entire filtered field, greatly reducing the computational cost. The filtering is performed for nine different averaging lengths: $L=3,6,12,24,48,96,192,384,$ and $768$ m.
@@
Calculating the drag force (16) also requires care as the line integral to evaluate since the building surfaces contain many horizontal surfaces for which $|\boldsymbol{N}_{\perp}|=0$. In order to ensure an appropriate discretisation of this term, note that by definition, for any boundary quantity $\phi$,
$\int\oint_{\partial\Omega_{f}(z)}\frac{\phi}{|\boldsymbol{N}_{\perp}|}\textnormal{d}s\textnormal{d}z=\iint_{\partial\Omega_{f}}\phi\textnormal{d}S,$ (29)
@@
$\sum_{k}\smash{\left(\sum_{m:K_{m}=k}\frac{\phi_{m}}{|\boldsymbol{N}_{\perp}|_{m}}\Delta s_{m}\right)}\Delta z_{k}=\sum_{m}\phi_{m}A_{m},$ (30)
@@
$\frac{\Delta s_{m}}{|\boldsymbol{N}_{\perp}|_{m}}=\frac{A_{m}}{\Delta z_{K_{m}}},$ (31)
which clarifies how ${\Delta s_{m}}/{|\boldsymbol{N}_{\perp}|_{m}}$ needs to be evaluated; it also shows that this term is always nonzero and finite. In order to be able to straightforwardly incorporate filtering, the surface quantity $\phi$ is converted to a surface density field $\rho_{\phi}$ that contains the boundary term and which is only non-zero in the cells next to the boundary:
@@
where $M_{ijk}\in\{m:I_{m}=i,J_{m}=j,K_{m}=k\}$ is the set of all cell-facets associated with cell $i,j,k$. Using (32), any surface quantity can be filtered straightforwardly using $\left\langle{\smash{\rho_{\phi}}}\right\rangle(\boldsymbol{x})=\int_{\Omega_{f}}\mathcal{A}(\boldsymbol{x}-\boldsymbol{y})\rho_{\phi}(\boldsymbol{y})\textnormal{d}\boldsymbol{y}$. Here we introduce a volumetric field for two surface quantities:

### v2

@@ 2.3 Triple decomposition
We define the Reynolds average of $\varphi$ as an ensemble average (Schmid et al. 2019):
$\overline{\varphi}(\boldsymbol{x},t)\equiv\sum_{i=1}^{N}\varphi^{(i)}(\boldsymbol{x},t)\,.$ (10)
Deviations from the average value $\overline{\varphi}$ are denoted $\varphi^{\prime}(\boldsymbol{x},t)=\varphi(\boldsymbol{x},t)-\overline{\varphi}(\boldsymbol{x},t)$. For statistically steady problems, time-averaging can be used in lieu of an ensemble average. In the current paper, we will assume that Reynolds-averaging has been performed before spatial averaging, which avoids some of the complexities of simultaneous Reynolds and spatial averaging (Schmid et al. 2019). We thus decompose the average velocity into the local spatial average and its deviation from the average as $\overline{\varphi}=\left\langle{\smash{\overline{\varphi}}}\right\rangle{{}^{f}}+\overline{\varphi}^{\prime\prime}$. This decomposition is only defined in the fluid domain, and the intrinsic average is used since this represents the average value inside the fluid domain. Substitution into Eq. (10) results in the triple decomposition
@@
For products, e.g. of a scalar flux $u_{i}\varphi$, this implies
[…]
@@ 2.4 Reynolds- and area-averaged streamwise momentum equations
where $\delta_{ij}$ is the Kronecker delta. Superficial averaging and application of the spatial averaging theorem Eq. (7) results in
@@
$\left\langle{\smash{f_{D}}}\right\rangle=-\oint_{\partial\Omega_{f}}\mathcal{A}\frac{F_{i}N_{i}}{|\boldsymbol{N}_{h}|}\textnormal{d}s=-\oint_{\partial\Omega_{f}}\mathcal{A}\,\left(\overline{p}\frac{N_{x}}{|\boldsymbol{N}_{h}|}-\nu\frac{\partial\overline{u}}{\partial x_{j}}\frac{N_{j}}{|\boldsymbol{N}_{h}|}\right)\textnormal{d}s\,.$ (16)
This equation explicitly provides the drag force induced by the building surfaces that is comprised of a form drag associated with the pressure and a skin drag associated with the viscosity (Raupach and Shaw 1982; Finnigan 2000; Nepf 2012; Sützl et al. 2020).
Considering the fact that NWP models take into account the net effect of buildings without actually resolving them, the governing equations are posed in terms of superficial quantities. Recasting Eq. (15) in the standard form for the streamwise momentum equation, we thus obtain:
@@
$\left\langle{\smash{\tau_{i}}}\right\rangle=\left\langle{\smash{\overline{u}_{i}}}\right\rangle\left\langle{\smash{\overline{u}}}\right\rangle-\left\langle{\smash{\overline{u}_{i}\overline{u}}}\right\rangle-\left\langle{\smash{\overline{u^{\prime}_{i}u^{\prime}}}}\right\rangle\,$, (18)
is the unresolved kinematic stress term. Note that we have not decomposed the term $\left\langle{\smash{\overline{u}_{i}\overline{u}}}\right\rangle$ into its components and that the viscous flux has been omitted since it is negligible in magnitude compared to the other fluxes. Equation (17) highlights the two terms that need to be closed in order to represent the effect of buildings in a non-building resolving NWP model: 1) the distributed drag contribution $f_{D}$; and 2) the unresolved kinematic stress $\tau_{i}$.
@@ 2.5 Statistically steady flow with periodic boundary conditions
The case considered in this paper is a turbulent urban boundary layer flow with periodic boundary conditions and a free-slip top boundary at $z=h$ in a statistically steady state. In this situation, Eq. (17) becomes
@@
which is not much of a simplification. No term except the time derivative can be assumed zero for an arbitrary averaging length $L$. However, if the averaging length scale is very large, the problem will become homogeneous in $x$ and $y$ and all the horizontal derivatives will vanish. With this in mind, it is useful to rewrite the equation above as
$\frac{\partial\left\langle{\smash{\overline{u}}}\right\rangle\left\langle{\smash{\overline{u}}}\right\rangle}{\partial x}+\frac{\partial\left\langle{\smash{\overline{v}}}\right\rangle\left\langle{\smash{\overline{u}}}\right\rangle}{\partial y}+\frac{\partial\left\langle{\smash{\overline{p}}}\right\rangle}{\partial x}-\frac{\partial\left\langle{\smash{\tau_{x}}}\right\rangle}{\partial x}-\frac{\partial\left\langle{\smash{\tau_{y}}}\right\rangle}{\partial y}=\left\langle{\smash{f}}\right\rangle-\left\langle{\smash{f_{D}}}\right\rangle-\frac{\partial\left\langle{\smash{\overline{w}}}\right\rangle\left\langle{\smash{\overline{u}}}\right\rangle}{\partial z}+\frac{\partial\left\langle{\smash{\tau_{z}}}\right\rangle}{\partial z}\,$, (20)
Here, the horizontal terms on the left-hand side will vanish if the problem is homogeneous, leaving the terms on the right-hand side of the equation to balance. The equation above indicates that an imbalance in the terms on the right-hand side implies inhomogeneity in $x$ and $y$ via horizontal terms on the left-hand side and vice versa.
@@
$\left\langle{\smash{\tau_{z}}}\right\rangle_{\infty}=-\left\langle{\smash{\overline{w}\,\overline{u}}}\right\rangle_{\infty}-\left\langle{\smash{\overline{w^{\prime}u^{\prime}}}}\right\rangle_{\infty}=-\left\langle{\smash{\overline{w}^{\prime\prime}\overline{u}^{\prime\prime}}}\right\rangle_{\infty}-\left\langle{\smash{\overline{w^{\prime}u^{\prime}}}}\right\rangle_{\infty}\,$, (22)
note that $\left\langle{\smash{\overline{w}}}\right\rangle_{\infty}=0$ due to volume conservation. The subscript on the superficial average indicates the averaging length $L$, in this case $L/\ell\gg 1$ which we denote as $L=\infty$. In the limit of $L\gg\ell$, the quantity $\left\langle{\smash{\varphi}}\right\rangle_{\infty}$ is identical to the conventional plane-average $\left\langle{\smash{\varphi}}\right\rangle_{xy}$, which is typically used in computational studies (Coceal et al. 2006; Xie et al. 2008; Giometto et al. 2016; Nazarian et al. 2020; Sützl et al. 2020). The superficial plane average of $\varphi$ is defined as
$\left\langle{\smash{\varphi}}\right\rangle_{xy}(z,t)\equiv\frac{1}{A_{T}}\int_{\Omega_{f}(z)}\varphi(\boldsymbol{x},t)\,\textnormal{d}\boldsymbol{x}_{h}\,$, (23)
@@
Integrating Eq. (21) from height $z$ to the domain top at height $h$ results in the following cumulative stress relation (Sützl et al. 2020):
@@
$\tau_{f;L}(\boldsymbol{x}_{h},z)$ $\equiv\int_{z}^{h}\left\langle{\smash{f}}\right\rangle_{L}(\boldsymbol{x}_{h},z^{\prime})\textnormal{d}z^{\prime}\,$,
$\tau_{D;L}(\boldsymbol{x}_{h},z)$ $\equiv\int_{z}^{h}\left\langle{\smash{f_{D}}}\right\rangle_{L}(\boldsymbol{x}_{h},z^{\prime})\textnormal{d}z^{\prime}\,$,
@@
In order to keep consistency in notation with earlier work, we will use that in the absence of a specific filter size in the subscript of a quantity, it is implied that $L=\infty$, e.g. $\tau_{f}\equiv\tau_{f;\infty}$; and we will use the conventional plane average notation $\left\langle{\smash{\varphi}}\right\rangle_{xy}$ instead of its equivalent $\left\langle{\smash{\varphi}}\right\rangle_{\infty}$. With this in mind, the cumulative stress relation Eq. (24) is simply
@@ 3.1 Case setup
In order to investigate multi-scale aspects of urban flows, we consider an idealised quasi-realistic urban morphology that has a staggered arrangement of cuboidal buildings of breadth and width $W=24\,\textnormal{m}$, whose height $H$ is random (Fig. 2). This case is inspired by Xie et al. (2008), but with a much larger simulation domain of $L_{x}\times L_{y}\times L_{z}=1\,536\,\textnormal{m}\times 768\,\textnormal{m}\times 512\,\textnormal{m}$, and using a height distribution representative for London (Sützl et al. 2021) rather than a Gaussian height distribution.
@@
The mean building height is set to $h_{m}=30\,\textnormal{m}$ and the ratio is set to $r=2.5$, both of which are close to the average value of London reported by Sützl et al. (2021). Thus the maximum building height is $h_{\max}=75\,\textnormal{m}$ and $\alpha=2.61$. At $z=0$, the total building width $L_{b}(z;\alpha)=N_{b}W$, which implies that the frontal area can be determined from Eq. (26) as
@@
In order to assign heights to each cuboidal building, note that Eq. (26) divided by $W$ can be inverted to show the height $z$ as a function of the number of buildings $L_{b}(z)/W$, which represents the distribution of heights as a function of the number of buildings (Fig. 2(a), solid line). The building heights can be determined by uniformly sampling this function for each $i=1,2,\ldots,N_{b}$. The first X buildings, where the function is not defined, are set to $h_{\max}$, and the minimum building height is set to $10$ m. As a final step, the heights are randomly distributed using a uniform distribution to all $512$ buildings as shown in the morphology in Fig. 2(b). The total building width $L_{b}$ is shown in Fig. 6(a) against the height, above $z=10$ m which is the minimum building height, the building width $L_{b}$, also indicating that the distributive frontal area smoothly decreases to zero at the top of the canopy region. Due to the minimum building height setting, the actual frontal area index is $\lambda_{f}=0.29$, slightly greater than previously designed.
@@
The urban flows over this morphology are performed by the large-eddy simulation using our open-source code uDALES (Owens et al. 2024), presenting the solid boundary using the immersed boundary method (IBM). The dynamics near the boundary are parametrised by the logarithmic wall functions (Uno et al. 1995; Suter et al. 2022). The turbulent eddy viscosity is calculated following the Vreman (2004) subgrid model. The code employs a second-order central difference scheme on a staggered Arakawa C-grid for spatial discretisation and an explicit third-order Runge-Kutta scheme for time integration.
The simulation is neutral and a constant pressure gradient forcing $\textnormal{d}P/\textnormal{d}x=0.0042$ kg m-2s-2 is imposed on the streamwise ($x$) direction to drive the wind. The simulation has a grid number of $N_{x}\times N_{y}\times N_{z}=1\,024\times 512\times 512$ (i.e., the grid size is $\Delta x\times\Delta y\times\Delta z=1.5\,\textnormal{m}\times 1.5\,\textnormal{m}\times 1\,\textnormal{m}$), with periodic boundary conditions applied in the lateral sides while the domain top is free-slip. The simulation runs for 10 000 s and the last 8 000 s is used to obtain converged time-averaged statistics.
3.2 Implementation details
@@
Efficient computation of the convolution integral Eq. (2) is crucial to carry out a multi-scale analysis. In the spatial domain, the convolution integral needs to be evaluated $N_{x}\times N_{y}\times N_{z}$ times for a single averaging length $L$ which is computationally very expensive, particularly for high-resolution simulations: assuming a filter contains $n^{2}$ grid cells, the convolution over the entire domain requires the number of operations of $O(n^{2}N_{x}^{2}N_{z})$ where we have assumed $N_{x}=N_{y}$. This is acceptable for small filter scales but becomes computationally inhibiting for large filter sizes. Therefore, we perform the convolution in the Fourier space by making use of the convolution theorem, which states that a convolution becomes a multiplication in Fourier space (Bracewell 2000), thus reducing the computational complexity of a 1-D convolution to $O(N_{x}\log N_{x})$. We thus compute the horizontal 2-D FFT of $I_{f}\varphi$ and the filter $\mathcal{A}$, multiply them, and then transform the product back to physical space using an inverse FFT, obtaining the filtered results for all points in the plane simultaneously. Repeating this for all $N_{z}$ planes requires $O((\log N_{x})^{2}N_{x}^{2}N_{z})$ operations, thus greatly reducing the computational cost, particularly for large simulation domains and large filters. The algorithm can be downloaded from Van Reeuwijk and Huang (2025). The filtering is performed for nine different averaging lengths: $L=3,6,12,24,48,96,192,384,$ and $768$ m.
@@
Calculating the drag force Eq. (16) also requires care as the line integral to evaluate since the building surfaces contain many horizontal surfaces for which $|\boldsymbol{N}_{h}|=0$. In order to ensure an appropriate discretisation of this term, note that by definition, for any boundary quantity $\phi$,
$\int\oint_{\partial\Omega_{f}(z)}\frac{\phi}{|\boldsymbol{N}_{h}|}\textnormal{d}s\textnormal{d}z=\oiint_{\partial\Omega_{f}}\phi\textnormal{d}S,$ (29)
@@
$\sum_{k}\smash{\left(\sum_{m:K_{m}=k}\frac{\phi_{m}}{|\boldsymbol{N}_{h}|_{m}}\Delta s_{m}\right)}\Delta z_{k}=\sum_{m}\phi_{m}A_{m},$ (30)
@@
$\frac{\Delta s_{m}}{|\boldsymbol{N}_{h}|_{m}}=\frac{A_{m}}{\Delta z_{K_{m}}},$ (31)
which clarifies how ${\Delta s_{m}}/{|\boldsymbol{N}_{h}|_{m}}$ needs to be evaluated; it also shows that this term is always nonzero and finite. In order to be able to straightforwardly incorporate filtering, the surface quantity $\phi$ is converted to a surface density field $\rho_{\phi}$ that contains the boundary term and which is only non-zero in the cells next to the boundary:
@@
where $M_{ijk}\in\{m:I_{m}=i,J_{m}=j,K_{m}=k\}$ is the set of all cell-facets associated with cell $i,j,k$. Using Eq. (32), any surface quantity can be filtered straightforwardly using $\left\langle{\smash{\rho_{\phi}}}\right\rangle(\boldsymbol{x})=\int_{\Omega_{f}}\mathcal{A}(\boldsymbol{x}_{h}-\boldsymbol{y}_{h})\rho_{\phi}(\boldsymbol{y}_{h})\textnormal{d}\boldsymbol{y}_{h}$. Here we introduce a volumetric field for two surface quantities: